# Survival Analysis with Adversarial Regularization

### v4

@@ Abstract
Survival Analysis (SA) models the time until an event occurs, with applications in fields like medicine, defense, finance, and aerospace. Recent research indicates that Neural Networks (NNs) can effectively capture complex data patterns in SA, whereas simple generalized linear models often fall short in this regard. However, dataset uncertainties (e.g., noisy measurements, human error) can degrade NN model performance. To address this, we leverage advances in NN verification to develop training objectives for robust, fully-parametric SA models. Specifically, we propose an adversarially robust loss function based on a Min-Max optimization problem. We employ CROWN-Interval Bound Propagation (CROWN-IBP) to tackle the computational challenges inherent in solving this Min-Max problem. Evaluated over 10 SurvSet datasets, our method, Survival Analysis with Adversarial Regularization (SAWAR), consistently outperforms baseline adversarial training methods and state-of-the-art (SOTA) deep SA models across various covariate perturbations with respect to Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI) metrics. Thus, we demonstrate that adversarial robustness enhances SA predictive performance and calibration, mitigating data uncertainty and improving generalization across diverse datasets by up to 150% compared to baselines.
Code: https://github.com/mlpotter/SAWAR
@@
A major challenge in SA is developing models that effectively handle complex time-to-event data while also ensuring robustness against dataset uncertainty, accurate calibration, and strong generalization. For instance, data uncertainty in medical applications can arise from various sources, including human factors such as illegible handwriting by physicians [7] and poor medical record-keeping [8], as well as inherent noise from medical equipment during data collection [9], termed as aleatoric uncertainty. Most of the existing approaches address aleatoric uncertainty by assuming noise in the logarithm of measured time-to-event data [10]. However, they often overlook uncertainty originating from the model’s input covariates.
SA models ought to generalize over various input covariates to sustain high accuracy in real-life applications. In medical studies, this entails high accuracy over diverse patient demographics, institutions, and data acquisition settings. While traditional SA models, such as Linear Cox-Weibull model [5], offer inductive biases specific to time-to-event analysis that aid generalization, they struggle with complex data relationships. On the other hand, while NN models can manage complex data, they may not generalize well to unseen data [11]. Combining the inductive biases of traditional SA models with NN techniques—such as modeling intensity or hazard functions with NNs —can both improve generalization and handle complex data [12]. Nonetheless, NN in SA may still suffer from poor model calibration, even with good generalization.
@@
Despite these challenges and the sensitivity of NNs to input uncertainties, robustness techniques for handling data perturbations in NN-based SA remain underexplored. To address these challenges, we make the following contributions:
We propose a novel training objective, called Survival Analysis with Adversarial Regularization (SAWAR), designed to create robust, fully-parametric NN-based SA models by incorporating adversarial covariate perturbations via convex relaxations, and thereby improve model calibration and generalization against noisy data.
Extensive experiments across 10 publicly-available SA datasets reveal the sensitivity of existing NN-based SA models to covariate perturbations with respect to Concordance Index (CI), Integrated Brier Score (IBS), and Negative Log Likelihood (NegLL) metrics, as well as survival curves.
SAWAR enhances SA predictive performance and calibration, mitigating data uncertainty and improving generalization by up to 150% compared to baseline adversarial training methods and state-of-the-art (SOTA) NN-based SA models.
@@ III Preliminaries on Survival Analysis
SA [21] is a statistical approach used to analyze and model the time until an event happens (e.g., a disease occurs, a device breaks). Typically, a study of such events happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow-up. In those cases, the dataset records that no event occurred up until the exit time of the study, which is called right censoring [45]. Such an SA dataset consists of independent instances, each denoted by index , for which the time between entry to a study and the subsequent event is recorded. Upon entry time to a study, each instance’s covariates (i.e., variables that govern the event of interest) are measured. A visual of time-to-event data with right censoring is shown in Fig. 1. In summary, the dataset has the following structure:
Being a statistical approach, we define the required probabilistic functions to model time-to-event data next.
@@ III-A Cox Proportional Hazard Model
Let us denote as a continuous, non-negative, random variable that represents the time-to-event, with Probability Density Function (PDF) and Cumulative Distribution Function (CDF) . The survival function is the complement of this CDF [46, 21].
An exponential SA model has the parametric form:
@@
is the instantaneous failure rate at time .
To incorporate each instance’s covariates in predicting failure rate, we use the prevalent CPH model [5]:
where is a baseline hazard, is the relative hazard associated with covariates , and is an NN with parameters in our setting.
The exponential CPH model assumes a time-independent baseline hazard rate . This allows us to “absorb” the term into the bias of the NN:
Further, the exponential CPH model’s PDF, CDF, and complement CDF are given by , , , respectively. is also known as the instance survival function.
The population survival curve is defined as the marginalization of the instance survival function over the covariates:
Since the distribution of covariates is typically unknown, the integral in Eq. 6 is often replaced with a Monte Carlo estimate over a dataset [31]:
Given these definitions, the next section describes the baseline objective for estimating model parameters from dataset .
@@ III-B Baseline Objective Functions
III-B1 Right Censored Log Likelihood Objective
@@
The first summation in Eq. 8 is the log likelihood of the datapoints corresponding to when the event occurs and the exact time that occurrence is observed. The second summation in Eq. 8 is the log likelihood when right censoring occurs, where it is known that no event occurred at least up until the censoring time.
@@ III-B2 Ranking Objective
The objective is a ranking loss that penalizes an incorrect ordering of pairs , where instance with event time should have a higher probability of failure than instance at time :
where and indicates an “acceptable” comparison pair [47, 48]. Two instances are comparable if the instance with lower observed time experienced an event, i.e., if . We set for all experiments. This objective is a continuous differentiable proxy for the non-continuous non-differentiable concordance index [49].
@@ III-B3 Combined Objective
The combined objective over the right censored dataset is:
where is the hazard versus ranking objective weight trade-off. Accordingly, is estimated by solving the following optimization problem:
This combined objective provides a trade-off between maximizing the likelihood of observing and maximizing a rank correlation between the instance hazard scores and the observed failure times. From here on, we term this combined loss as the baseline objective. We note that an NN-based SA model DRAFT [25] also employs a combined objective. Our baseline objective incorporates a different ranking objective than DRAFT and outperforms DRAFT as discussed in Section VI.
@@ IV Preliminaries on Adversarial Training
Adversarial training [36, 50, 51] improves robustness and generalization of NNs to uncertain data. Rather than training on pristine data, most adversarial training approaches perturb the data throughout the training process to encourage the model to optimize for parameters that are resilient against such perturbations. For example, [13] formulated adversarial training as solving a Min-Max robust optimization problem:
@@
In practice, the Min-Max problem posed in [13] is intractable to solve exactly. In particular, the inner maximization requires finding a global optimum over a high dimensional and non-convex loss function over input covariates with the current NN parameters. Therefore, many recent adversarial training methods propose approximations to Eq. 12, leading to models with varying degrees of robustness to different data uncertainties. To understand the robustness properties of SA models to a variety of data uncertainties present in real applications, this paper investigates several perturbation methods, described next.
@@ IV-A Verifiable Robustness
One such approximation is to use techniques from NN verification during training. Neural network robustness verification algorithms determine the output neuron’s upper and lower bounds for all inputs in a set , typically constrained by a norm-bounded perturbation [51]. For example, CROWN-Interval Bound Propagation (CROWN-IBP) combines CROWN and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data within the set . [50, 51, 52]. CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs. [53, 50] showed the linear relaxation can be extended to the entire objective function:
where is a convex relaxation upper bound of . This relaxation is convex with respect to NN weights, enabling efficient optimization, and by optimizing over an upper bound on the inner maximization, it accounts for a strong adversary during training.
@@ IV-B Adversarial Perturbations
Another way to approximate Eq. 12 is to use adversarial perturbations.
@@ IV-B1 Fast Gradient Sign Method (FGSM)
FGSM [54] uses the gradient of the objective function with respect to the input to perturb in the direction that maximizes the objective . Projected Gradient Descent (PGD) [13] iteratively applies that same type of first-order perturbation times (or until convergence is reached), but constrains the perturbation of to remain within a specified set (e.g., -ball). FGSM is a special case of PGD, with and the -ball.
Considering the structure of the datasets as described in Section III, we perturb only the covariates and not the time-to-event . The perturbed covariates are calculated as:
@@ IV-B2 Random Noise
While robustness to a worst-case perturbation is important for many applications, adding random noise to training data has also been shown to improve generalization performance [55]. Therefore, this paper also considers Gaussian Noise perturbations with variance bounded within an -ball:
Building on this foundation, we next introduce our approach, SAWAR, that develops robust SA models based on adversarial regularization.
@@ V Survival Analysis with Adversarial Regularization
This paper addresses the open challenge in SA of developing models robust to dataset uncertainty, generalizable, and calibrated. We propose a training objective for adversarial robustness in fully-parametric NN-based SA, enhancing calibration and generalization on unexpected or noisy data. Our approach extends the objective function (Eq. 10) with adversarial regularization as a Min-Max optimization problem (Eq. 12) [51, 50], as detailed in the following sections.
@@ V-A Min-Max Formulation
When solving for in SA, many approaches treat each measurement in dataset as ground-truth, ignoring aleatoric uncertainty [10]. Typically in SA, the aleatoric uncertainty is only associated with the noise in the logarithm of measured time-to-event (e.g. the Gumbel distribution or Log Normal distribution) [21]. Instead, the proposed SAWAR approach assumes that each covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV. Thus, we optimize the NN parameters for the worst-case realizations of Eq. 10:
@@ VI Experiments
We conduct experiments over 10 benchmark medical SA datasets (Section VI-A) to quantify the impact of perturbation after the robust training is complete. To demonstrate that SAWAR produces SA models that are both high-performing and robust to dataset uncertainty, this section quantifies improvements in predictive accuracy, calibration, and population curve estimation compared to common adversarial training techniques (Section IV-B) including FGSM, Random Noise and PGD, as well as SOTA NN-based SA models, such as DRAFT and AAE-Cox.
@@ VI-B Experiment Setup
As evaluation metrics, we consider the CI [49], the IBS [69] and the NegLL at varying -perturbation magnitudes from 0 to 1. We use lifelines [70] and sksurv [48] for computing SA metrics.
@@ VI-C Adversary Setup
We apply two adversarial perturbation settings on the test set to assess the robustness of the proposed adversarial regularization method: FGSM perturbation (Section IV-B1) and a novel worst-case perturbation. The perturbed covariates with respect to the input are found via Eq. 15 and the corresponding failure rate becomes:
We define the worst-case perturbation as the maximum failure rate with respect to the covariate uncertainty:
which we solve via CROWN-IBP (Section IV-B1) by optimizing a tight approximation of the upper bound of the maximization term. Accordingly, the -worst-case survival curve has the lowest survival probability at a given time with respect to the -perturbed input covariates:
@@
For the exponential CPH distribution, as the failure rate increases with the -perturbation, the probability of survival decreases monotonically. We evaluate SA metrics for , where represents no perturbation to input covariates, in order to study how sensitive different adversarial training methods are to covariate perturbations.
To simulate realistic conditions, we assume that the 10 benchmark survival analysis datasets are “clean” and contain various covariates (e.g., age, drug dosage in mg, tumor size) and survival outcomes (e.g., time until death, censored observations). By applying adversarial attack techniques, we perturb these input covariates to maximize the survival analysis error.
The perturbations are designed to be subtle yet impactful, meaning they are small in magnitude but cause significant changes in the predicted survival times. For example, since the data is standard normalized, a perturbation of would shift an input covariate by 0.1 standard deviations from its current value.
For instance, consider a scenario where one of the covariates is the body weight of a patient in kilograms. If patient has a body weight of 70 kg, an adversarial perturbation might increase this body weight to to 75 kg. While this change is minimal, it is selected because it may cause a drastic change in the predicted survival likelihood over the next 10 years for a non-robust NN-based SA model. In reality, a 5 kg change in weight should not have such a significant effect on survival, but the model may incorrectly interpret this small perturbation as a substantial change in survival hazard, highlighting its sensitivity to minor changes in input covariates.

### v5

@@ Abstract
Survival Analysis (SA) models the time until an event occurs, with applications in fields like medicine, defense, finance, and aerospace. Recent research indicates that Neural Networks (NNs) can effectively capture complex data patterns in SA, whereas simple generalized linear models often fall short in this regard. However, dataset uncertainties (e.g., noisy measurements, human error) can degrade NN model performance. To address this, we leverage advances in NN verification to develop training objectives for robust, fully-parametric SA models. Specifically, we propose an adversarially robust loss function based on a Min-Max optimization problem. We employ CROWN-Interval Bound Propagation (CROWN-IBP) to tackle the computational challenges inherent in solving this Min-Max problem. Evaluated on 10 SurvSet datasets, Survival Analysis with Adversarial Regularization (SAWAR) consistently outperforms baseline adversarial training methods and state-of-the-art (SOTA) deep SA models on negative log-likelihood (Negative Log Likelihood (NegLL)), integrated Brier score (Integrated Brier Score (IBS)), and concordance index (Concordance Index (CI)). These improvements persist across a range of input covariate perturbations. Thus, we demonstrate that adversarial robustness enhances SA predictive performance and calibration, mitigating input covariate uncertainty and improving generalization across diverse datasets by up to 150% compared to baselines.
https://github.com/mlpotter/SAWAR
@@
A major challenge in SA is developing models that effectively handle complex time-to-event data while also ensuring robustness against dataset uncertainty, accurate calibration, and strong generalization. For instance, input covariate uncertainty in medical applications can arise from various sources, including human factors such as illegible handwriting by physicians [7] and poor medical record-keeping [8], as well as inherent noise from medical equipment during data collection [9], termed as aleatoric uncertainty. Most of the existing approaches address aleatoric uncertainty by assuming noise in the logarithm of measured time-to-event data [10]. However, they often overlook uncertainty originating from the model’s input covariates.
SA models should generalize over various input covariates to maintain high accuracy in real-life applications. In medical studies, this entails high accuracy over diverse patient demographics, institutions, and data acquisition settings. While traditional SA models, such as Linear Cox-Weibull model [5], offer inductive biases specific to time-to-event analysis that aid generalization, they struggle with complex data relationships. On the other hand, while NN models can manage complex data, they may not generalize well to unseen data [11]. Combining the inductive biases of traditional SA models with NN techniques, such as modeling intensity or hazard functions with NNs, can both improve generalization and handle complex data [12]. Nonetheless, NNs in SA may still suffer from poor model calibration, even with good generalization.
@@
Despite these challenges and the sensitivity of NNs to input uncertainties, robustness techniques for handling input covariate perturbations in NN-based SA remain underexplored. To address these challenges, we make the following contributions:
We propose a novel training objective, called Survival Analysis with Adversarial Regularization (SAWAR), designed to create robust, fully-parametric NN-based SA models. We achieve this by adding an adversarial regularization term formulated as a Min-Max objective. The inner maximization is approximated with convex relaxations (e.g., CROWN-IBP), producing a tractable upper bound used during training to improve calibration and generalization under covariate uncertainty.
Extensive experiments on 10 publicly available SA datasets demonstrate that current adversarial training methods for NN-based SA models are highly sensitive to covariate perturbations. We quantify this sensitivity using Concordance Index (CI), Integrated Brier Score (IBS), and Negative Log Likelihood (NegLL) metrics and by comparing the resulting population survival curves under various levels of input covariate perturbation.
SAWAR enhances SA predictive performance and calibration, mitigating input covariate uncertainty and improving generalization by up to 150% compared to baseline adversarial training methods and state-of-the-art (SOTA) NN-based SA models.
@@ III Preliminaries on Survival Analysis
SA [21] is a statistical approach used to analyze and model the time until an event occurs (e.g., a disease onset or device failure). Instances/individuals may leave a study before an event is observed; this is recorded as right censoring [45]. Each of the independent instances/individuals (indexed by has: covariates measured at study entry, a recorded time , and an event indicator . We use the convention:
: the event was observed for instance ; then is the event time.
: instance is right-censored; then is the censoring time (the last time the instance was known to be event-free).
A visual of time-to-event data with right censoring is shown in Fig. 1. In summary, the dataset is written as
We next define the required probabilistic functions to model time-to-event data next.
@@ III-A Cox Proportional Hazard Model
Let us denote as a continuous, non-negative, random variable that represents the time-to-event, with Probability Density Function (PDF) and Cumulative Distribution Function (CDF) . The survival function is the complement of this CDF [46, 21]. A SA model has the parametric form:
@@
is the instantaneous risk at time . Informally, is the probability that an event occurs within a small time window , given that the event has not yet occurred.
To incorporate each instance’s covariates in predicting hazard, we use the prevalent CPH SA model formulation [5]:
where is a baseline hazard, is the relative hazard (or relative risk) associated with covariates , and is an NN with parameters in our setting. The exponential CPH model assumes a time-independent baseline hazard rate , which allows us to “absorb” the term into the bias of the NN:
Further, the exponential CPH model’s PDF, CDF, and complement CDF are given by , , , respectively. is also known as the instance survival function. The population survival curve is defined as the marginalization of the instance survival function over the input covariates:
Since the distribution of input covariates is typically unknown, the integral in Eq. 6 is often replaced with a Monte Carlo estimate over a dataset [31]:
Given these definitions, the next section describes the baseline objective for estimating SA model parameters from dataset .
@@ III-B Baseline Objective Functions
III-B1 Right-Censored Log Likelihood Objective
@@
The first summation in Eq. 8 is the log likelihood of the datapoints corresponding to when the event occurs and the exact time that occurrence is observed. The second summation in Eq. 8 is the log likelihood when right censoring occurs, where it is known that no event occurred at least up until the censoring time. A higher right-censored log likelihood indicates a better SA model fit to the data .
@@ III-B2 Ranking Objective
The objective is a ranking loss that penalizes an incorrect ordering of pairs , where instance with event time should have a higher probability of event occurrence than instance at time (:
where and indicates an “acceptable” comparison pair [47, 48]. Two instances are comparable if the instance with lower time experienced an event, i.e., if . We set for all experiments. This objective is a continuous differentiable proxy for the non-continuous non-differentiable concordance index [49].
@@ III-B3 Combined Objective
The combined objective over the right-censored dataset is:
where controls a trade-off between maximizing the right-censored likelihood of observing and maximizing a rank correlation Accordingly, is estimated by solving the following optimization problem:
@@ IV Preliminaries on Adversarial Training
Adversarial training [36, 50, 51] improves robustness and generalization of NNs to uncertain input covariates. Rather than training on “clean” data, most adversarial training approaches perturb the input covariates throughout the training process to encourage the NN model to optimize for parameters that are resilient against such perturbations. For example, [13] formulated adversarial training as solving a Min-Max robust optimization problem:
@@
In practice, the Min-Max problem posed in Eq. 12 [13] is intractable to solve exactly. In particular, the inner maximization requires finding a global optimum over a high dimensional and non-convex loss function over input covariates with the current NN parameters. Therefore, many recent adversarial training methods propose approximations to Eq. 12, leading to models with varying degrees of robustness to different input covariate uncertainties. To understand the robustness properties of SA models to a variety of input covariate uncertainties present in real applications, this paper investigates several perturbation methods, described next.
@@ IV-A Verifiable Robustness
One such approximation to the Min-Max objective in Eq. 12 is to use techniques from NN verification during training. NN robustness verification algorithms determine the output neuron’s upper and lower bounds for all input covariates in a set , typically constrained by a norm-bounded perturbation [51]. For example, CROWN-Interval Bound Propagation (CROWN-IBP) combines CROWN and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input covariates within the set [50, 51, 52]. CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs. [53, 50] showed the linear relaxation can be extended to the entire objective function:
where is a convex relaxation that upper-bounds . This relaxation is convex with respect to the NN weights, enabling efficient optimization. By optimizing an upper bound on the inner maximization, the procedure accounts for a strong adversary during training.
@@ IV-B Adversarial Perturbations
Another way to approximate the inner maximization of Eq. 12 is to use adversarial perturbations to the input covariates.
@@ IV-B1 Fast Gradient Sign Method (FGSM)
FGSM [54] uses the gradient of the objective function with respect to the input to perturb in the direction that maximizes the objective . Projected Gradient Descent (PGD) [13] iteratively applies that same type of first-order perturbation times (or until convergence is reached), but constrains the perturbation of to remain within a specified set (e.g., -ball of ). FGSM is a special case of PGD, with and the -ball.
We only perturb the input covariates and not the time-to-event . The perturbed input covariates are calculated as:
@@ IV-B2 Random Noise
While robustness to a worst-case perturbation is important for many applications, adding random noise to training data has also been shown to improve generalization performance [55]. Therefore, this paper also considers Gaussian Noise perturbations with variance bounded within an -ball of :
Explicitly modeling input covariate uncertainty via a Min–Max formulation enables principled propagation of aleatoric uncertainty. Techniques such as CROWN-IBP provide formal analytical upper bounds for predefined perturbation sets, whereas gradient-based methods (FGSM, PGD) produce empirical, sensitivity-driven perturbations without formal guarantees. Approaches that do not specify uncertainty sets (e.g., GAN-based methods or AAE-Cox) rely on learned latent representations and therefore lack formal coverage of uncertainty and direct control over the strength of adversarial regularization.
Building on this foundation of incorportating aleatoric uncertainty via a Min-Max objective, we next introduce our approach, SAWAR, that develops robust SA models based on adversarial regularization.
@@ V Survival Analysis with Adversarial Regularization
This paper addresses the open challenge in SA of developing SA models robust to input covariate uncertainty, generalizable, and calibrated. We propose a training objective for adversarial robustness in fully-parametric NN-based SA, enhancing calibration and generalization on unexpected or noisy data. Our approach extends the objective function (Eq. 10) with a Min-Max adversarial regularization term (Eq. 12) [51, 50], as detailed in the following sections.
@@ V-A Min-Max Formulation
When solving for in SA, many approaches treat each measurement in dataset as ground-truth, ignoring aleatoric uncertainty [10]. Typically in SA, the aleatoric uncertainty is only associated with the noise in the logarithm of measured time-to-event (e.g. the Gumbel distribution or Log Normal distribution) [21]. Instead, the proposed SAWAR approach assumes that each input covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV. Thus, we add an adversarial regularization term which optimizes the NN parameters for the worst-case realizations of Eq. 10:
@@ V-B Adversarial Robustness Regularization
In initial experiments, we conducted a hyper parameter search over to assess its impact on the survival curves. At , the objective function reduces to the CROWN-IBP Min-Max formulation, resulting in narrower credible intervals. This is expected, as minimizing deviations in the survival curve due to input covariate perturbations may be achieved by reducing reliance on the covariates. Thus, may lead to over regularization hurting predictive performance of individual patients, but should still have a good estimate of the global population survival curve. At , the objective aligned with the standard log-likelihood in traditional SA for exponential-Cox proportional hazard models. Therefore, we may view as a hyperparameter which controls the tradeoff between adversarial regularization and predictive performance. We chose to equally balance adversarial regularization and predictive performance.
We expect that the model’s adversarial robustness and predictive calibration should increase when incorporating input covariate uncertainty in our SAWAR objective [18] via the adversarial regularization term in Eq. 19 and show SAWAR’s improved performance via extensive experiments in the next section.
@@ VI Experiments
We conduct experiments over 10 benchmark medical SA datasets (Section VI-A) to quantify the impact of input covariate perturbation after the robust training is complete.
To evaluate SAWAR, we train one neural network per adversarial training method. Standard approaches approximate the inner maximization in Eq. 18 using FGSM (Eq. 15), PGD, or Gaussian noise (Section IV-B2). We also include DRAFT and AAE-Cox as baselines, which alter the training objective or NN architecture. For each trained model, we then perturb test covariates by ( means no perturbation) and evaluate CI, IBS, and NegLL at each . Metrics are computed using lifelines [56] and sksurv [48].
@@ VI-C Adversary Setup
We apply two adversarial perturbation methods on the test set to assess the robustness of the proposed adversarial regularization method: FGSM perturbation (Section IV-B1) and a novel worst-case perturbation. The perturbed covariates with respect to the input are found via Eq. 15 and the corresponding hazard becomes:
We define the worst-case perturbation as the maximum hazard with respect to the input covariate uncertainty set:
which we solve via CROWN-IBP (Section IV-B1) by optimizing a tight approximation of the upper bound of the maximization term.
For the exponential CPH distribution, as the hazard increases with the -perturbation, the probability of survival decreases monotonically. Accordingly, the -worst-case survival curve has the lowest survival probability at a given time with respect to the -perturbed input covariates:
We evaluate SA metrics for , where represents no perturbation to input covariates, in order to study how sensitive different adversarial training methods are to covariate perturbations.
@@
To simulate realistic conditions, we assume that the 10 benchmark SA datasets are “clean” and contain various input covariates (e.g., age, drug dosage in mg, tumor size) and survival outcomes (e.g., time until death, censored observations). By applying adversarial attack techniques, we perturb these input covariates to maximize the SA error.
The perturbations are designed to be subtle yet impactful, meaning they are small in magnitude but cause significant changes in the predicted survival times. Since the data is standard normalized, a perturbation of would shift an input covariate by 0.1 standard deviations from its current value. For example, increasing a patient weight from 70 kg to 72 kg (a small real‑world change that would not alter clinical categorization of “Normal weight” according to National Institute of Health (NIH) [70]) may still produce a disproportionate change in the patient’s predicted survival for a sensitive model, illustrating the need for adversarial regularization.